# Boosting Graph Pooling with Persistent Homology

### v1

@@ 4.3 Analysis
Proposition 1. TIP is isomorphic invariant.
@@
Complexity. PH can be efficiently computed for dimensions 0 and 1, with a worst-case time complexity of $O(m\alpha(m))$, where $m$ represents the number of sorted edges in a graph. Here, $\alpha(\cdot)$ represents the inverse Ackermann function, which is extremely slow-growing and can essentially be considered as a constant for practical purposes. Therefore, the primary factor that affects the calculation of PH is the complexity of sorting all the edges, which is $O(m\log m)$. Our resampling and persistence encoding mechanism ensures that the coarsened graphs are sparse rather than dense, making our approach both efficient and scalable. Besides, we provide running time comparisons in Appendix E.2, which indicates that the inclusion of TIP does not impose a significant computational burden.
@@ 5.1 Experimental Setup
To investigate the effectiveness of PH in GP, we integrate TIP with DiffPool, MinCutPool, and DMoNPool, which are the pioneering approaches that have inspired many other pooling methods. Additionally, as most pooling methods rely on GNNs as their backbone, we compare the widely used GNN models GCN (Kipf & Welling, 2016), GIN (Xu et al., 2018), and GraphSage (Hamilton et al., 2017). We also look into another two related and State-of-the-Art GNN models, namely TOGL (Horn et al., 2021) and GSN (Bouritsas et al., 2022), which incorporate topological information and graph substructures into GNNs to enhance the expressive power. Furthermore, we compare several other GP methods, namely Graclus (Dhillon et al., 2007) and TopK (Gao & Ji, 2019). For model selection, we follow the guidelines provided by the original authors or benchmarking papers. Our method acts as an additional plug-in to existing pooling methods (referred to as -TIP) without modifying the remaining model structure and hyperparameters. Appendix B.1 provides detailed configurations of these models.
@@
To evaluate the capabilities of our model across diverse domains, we assess its performance on a variety of graph datasets commonly used in graph classification tasks. We select several benchmarks from TU datasets (Morris et al., 2020) and OGB datasets (Hu et al., 2020). Specifically, we adopt molecular datasets NCI1, NCI109, and OGBG-MOLHIV, bioinformatics datasets ENZYMES, PROTEINS, and DD, as well as social network datasets IMDB-BINARY and IMDB-MULTI. Furthermore, to investigate the topology-preserving ability of our method, we conduct experiments on several highly structured datasets (ring, torus, grid2d) obtained from the PyGSP library. Appendix B.2 provides detailed statistics of the datasets.
@@ 5.2 Preserving Topological Structure
In this experiment, we study Q1 about the ability of PH to preserve topological structure during pooling. Specifically, we assess the topological similarity between the original and coarsened graphs $\mathcal{G}$ and $\mathcal{G^{\prime}}$, by comparing the Wasserstein distance associated with their respective PDs $\mathcal{D}_{1}$ and $\mathcal{D^{\prime}}_{1}$. This evaluation criterion is widely used to compare the topological similarity of graphs (Yan et al., 2022; Southern et al., 2023). We utilize Forman curvature on each edge of the graph as the filtration, which incorporates edge weights and graph clusters to better capture the topological features of the coarsened graphs (Sreejith et al., 2016; Wee & Xia, 2021). We consider the 1-Wasserstein distance $W\left(\mathcal{D}_{1},\mathcal{D}^{\prime}_{1}\right)=\inf_{\gamma\in\Pi\left(\mathcal{D}_{1},\mathcal{D}^{\prime}_{1}\right)}\mathbb{E}_{(x,y)\sim\gamma}[\|x-y\|]$ as the evaluation metric, where $\Pi(\cdot)$ is the set of joint distributions $\gamma(x,y)$ whose marginals are $\mathcal{D}_{1}$ and $\mathcal{D}^{\prime}_{1}$, respectively. Note that we are not learning a new filtration but keep a fixed one. Rather, we use learnable filtrations in training to enhance flexibility, and solely optimize $L_{topo}$ as the objective.
Table 1: Results to show the topology-preserving ability of different pooling methods. We use the Wasserstein distance to assess the topological similarity. A bold value indicates the overall winner.
@@
Table 2: Test accuracy of graph classification on benchmark datasets. A bold value indicates the overall winner. Gray background indicates that TIP outperforms the base GP.
@@ 5.3 Graph Classification
In this experiment, we examine the impact of PH on GP in downstream tasks to answer Q2. We have observed in the former experiment that PH can preserve essential topological information during pool. However, two additional concerns arise: (1) Does TIP continue to generate invariant sub-topology in the downstream task? (2) If so, does this sub-topology contribute to the performance of the downstream task? To address these concerns, we evaluate TIP using various graph classification benchmarks, where the accuracy achieved on these benchmarks serves as a measure of a method’s ability to selectively preserve crucial information based on the task at hand.
We begin by visualizing the coarsened graphs in this task, where edges are cut-off by a small value. From Fig. 4, we can clearly observe that our method manage to preserve the essential sub-topology similar to the original graphs, while dense pooling methods cannot preserve any topology. As discussed in Mesquita et al. (2020), dense pooling methods achieve comparable performance when the assignment matrix $\mathbf{S}$ is replaced by a random matrix. Here our visualization reveals that regardless of the value of $\mathbf{S}$, the coarsened graph always approaches to a fully connected one. Sparse pooling methods, on the other hand, manage to preserve some local structures through clustering or dropping, but the essential global topological structures are destroyed.
Table 2 presents the average and standard deviation of the graph classification accuracy on benchmark datasets, where the results of GP and several baseline GNNs are provided. Experimental results demonstrate that TIP can consistently enhance the performance of the three dense pooling methods. While the original dense pooling methods sometimes underperform compared to the baselines, they are able to surpass them after integrating TIP. Among the different variants of dense pooling methods, DiffPool-TIP achieves the highest accuracy in most cases, this may be attributed to the fact that DiffPool applies three consecutive layers of GNNs after each pooling operation, while the other two methods only utilize one GNN layer. The coarsened graphs with invariant sub-topology are mostly sparsely connected, so additional layers of GNN may have a positive impact on passing crucial messages.
@@
Moreover, an intriguing observation can be found on ENZYMES dataset, where TOGL surpasses the baseline GNNs. TOGL in practice, incorporates PH into GNNs (GraphSage in our implementation), so this results underscores the significance of incorporating topological information for improved performance on ENZYMES. Further, our method demonstrates more significant improvements by augmenting the three dense pooling methods on the ENZYMES dataset. One possible explanation for the observed phenomenon is that the coarsened graphs generated by our methods bear a striking resemblance to numerous frequent subgraphs present in this dataset (Fu et al., 2023). Such substructures may serve as indicators of unique characteristics within the graph, rendering them valuable for subsequent tasks. However, it is also worth noting that TOGL only exhibits marginal improvements or even underperforms on the other datasets. This suggests that simply integrating PH features into GNN layers does not fully exploit topological information. Conversely, injecting global topological invariance into pooling layers in our method yields superior performance.
To demonstrate the effectiveness of preserving the invariant sub-topology, we compared DiffPool-TIP with its invariant counterpart, DiffPool-TIP-NL (no topological loss), by replacing $\mathcal{L}_{topo}$ with the original $\mathcal{L}_{r}$ in DiffPool (see Table 3 in Appendix A). The training objective curve and the Wasserstein distance curve are presented in Figure 5, both based on the ENZYMES dataset and a fixed filtration (the same as in Section 5.2). From the figures, it is evident that the objective value decreases as the coarsened graphs become more similar in topology to the original graphs when using DiffPool-TIP. However, when training without $\mathcal{L}_{topo}$, the performance is inferior. Additionally, even when the objective value converges, DiffPool-TIP-NL still exhibits changing topology, whereas DiffPool-TIP maintains a stable topology, possibly benefiting from the stability of PH (Southern et al., 2023). This also suggests that multiple suboptimal topologies may contribute equally to the objective. Our topology invariant pooling strategy consistently selects topologies similar to the original graph, which leads to better performance. Additional visualization results and analysis about the coarsened graphs obtained by DiffPool-TIP-NL can be found in Appendix E.3. Furthermore, ablation study about the contributions of different modules are shown in Appendix E.4.
Figure 5: The training curves of DiffPool-TIP and DiffPool-TIP-NL (no topological loss) on ENZYMES dataset. We show the average values and min-max range in terms of objective and Wasserstein distance for multiple runs.
@@ B.1 Implementation detail
In the graph classification task, all datasets are splitted into train (80%), validation (10%), and test (10%) data. Following the evaluation protocol in (Ying et al., 2018; Mesquita et al., 2020), we train all models using the Adam optimizer (Kingma & Ba, 2014) and implement a learning rate decay mechanism, reducing the learning rate from $10^{-3}$ to $10^{-5}$ with a decay ratio of 0.5 and a patience of 10 epochs. Additionally, we use early stopping based on the validation accuracy with patience of 50 epochs. We report statistics of the performance metrics over 20 runs with different seeds.
@@
All the methods are implemented using PyTorch and PyG (Paszke et al., 2017; Fey & Lenssen, 2019). The compared methods are implemented following the implementations provided in the PyG library  [^1] https://pytorch-geometric.readthedocs.io/en/latest/modules/nn.html. In the case of DiffPool, it uses a 3-layer GraphSage in each pooling layer, while MinCutPool and DMoNPool use a 1-layer GCN before pooling and a 1-layer GNN (Morris et al., 2019) in each pooling layer. Note that in DiffPool, the GNNs in Eqs. 2 and 4 are different, while in MinCut and DNoNPool they are the same one, as what their do in the original papers. TopK and Graclus are based on a 1-layer GNN (Morris et al., 2019). TOGL is implemented using a 3-layer GraphSage as it has demonstrated superior performance on graph classification tasks (see Table 2). For the baseline GNN models (GCN, GIN, and GraphSage), we use 3 layers with mean/max pooling. In our model, TIP is incorporated as a plugin to existing pooling methods, without modifying the remaining model structure and hyperparameters. We replace the reconstruction loss $\mathcal{L}_{r}$ with $\mathcal{L}_{topo}$ while keeping the cluster loss $\mathcal{L}_{c}$ unchanged. In the case of MinCutPool and DMoNPool, our resampling strategy is added after their original normalization of the coarsened graphs. In preserving topological structure experiments, we initialize node features as the concatenation of the first ten eigenvectors of graph Laplacian matrices.
@@ Appendix C Theoretical Expressivity of TIP
Proof. Assume that $\mathcal{G}$ and $\mathcal{G}^{\prime}$ have $n$ and $n^{\prime}$ nodes, and the label sequences of them diverge at some iteration $h$, which means there exists at least one label whose count is unique. Let nodes $u$ and $u^{\prime}$ be the nodes with unique count in $\mathcal{G}$ and $\mathcal{G}^{\prime}$, respectively. Denote $La^{(h)}:=\{l_{1},l_{2},...\}$ as an enumeration of the finitely many hashed labels at iteration $h$. We can build a filtration function $f$ by assigning a vertex $v$ with label $l_{i}$ to its index, i.e. $f(v):=i$ except that $f(u)=n+n^{\prime}+1$ and $f(u^{\prime})=n+n^{\prime}+2$. The filtration of edge $(u,w)$ is defined as $f(v,w):=max\{f(v),f(w)\}$, and for isolated nodes $v$, the filtration of self-loop edges is $f(v,v)=f(v)$. Therefore, node with unique label count and its connected edges always correspond to the largest filtration value. Note that the 1-dimensional PD has been extended to have the same cardinality as the number of edges. If node $u$ or $u^{\prime}$ forms a circle, the creation of this circle is related to the edge with the largest filtration; if node $u$ or $u^{\prime}$ does not form a circle, the corresponding edges lie on the diagonal of $\mathcal{D}_{1}$ with unique coordinates; otherwise node $u$ constitute a circle while $u^{\prime}$ does not, then the corresponding edges lie in different parts in $\mathcal{D}_{1}$ and $\mathcal{D}^{\prime}_{1}$. Hence, $\mathcal{D}_{1}\neq\mathcal{D}^{\prime}_{1}$.
@@
To demonstrate that TIP is more expressive than other dense pooling methods, we provide examples of graph pairs that cannot be distinguished by 1-WL but can be by TIP. We present an example of such non-isomorphic graphs in Fig. 6, where in the second graph the edge connecting two triangles does not form a circle. This edge corresponds to zero persistence and is eliminated in TIP. Consequently, the two originally non-isomorphic graphs can be easily distinguished. Provided that the three sufficient conditions proposed in (Bianchi & Lachi, 2023) are satisfied, the pooling layers retain the same level of expressive power as GNN. In TIP, the reduction of node features remains unaltered, thereby fulfilling the three conditions. Additionally, TIP is capable of distinguishing certain non-isomorphic graphs, indicating its superior expressive power compared to conventional dense pooling methods such as DiffPool, MinCutPool, and DMoNPool.
Proposition 1. TIP is isomorphic invariant.
@@
Lemma 1 (Theorem 2 in (Rieck, 2023)). Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be two isomorphic graphs. For any equivariant filtration $f$, the corresponding persistence diagrams are equal.
@@
Proof. Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be two isomorphic graphs with $n$ nodes. Their node features after applying several layers of GNNs are denoted as $\mathbf{X}_{1}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ and $\mathbf{X}_{2}=\{\mathbf{y}_{i}\}_{i=1}^{n}$, respectively. Let the assignment matrix of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ obtained in Eq. 2 be $\mathbf{S}_{1}=\left\{\left\{s_{ij}\right\}_{i=1}^{n}\right\}_{j=1}^{n^{\prime}}$ and $\mathbf{S}_{2}=\left\{\left\{q_{ij}\right\}_{i=1}^{n}\right\}_{j=1}^{n^{\prime}}$ with ${\textstyle\sum_{j=1}^{n^{\prime}}}s_{ij}=1$ and ${\textstyle\sum_{j=1}^{n^{\prime}}}q_{ij}=1$. Let the coarsened graphs after pooling as $\mathcal{G}_{1P}$ and $\mathcal{G}_{2P}$ and with $n^{\prime}$ nodes, and their corresponding node features as $\mathbf{X}_{P1}=\{\mathbf{x}_{Pj}\}_{j=1}^{n^{\prime}}$ and $\mathbf{X}_{P2}=\{\mathbf{y}_{Pj}\}_{j=1}^{n^{\prime}}$, where $x_{Pj}=\textstyle\sum_{i=1}^{n}\mathbf{x}_{i}\cdot\mathbf{s}_{ij}$ and $y_{Pj}=\textstyle\sum_{i=1}^{n}\mathbf{y}_{i}\cdot\mathbf{q}_{ij}$ according to Eq. 4.
@@
which indicates that the output of TIP is invariant given isomorphic graphs as input. We have already proved that isomorphic graph pairs after TIP are still isomorphic, so Eq. 9 still holds for multiple layers of pooling.

### v2

@@ 4.3 Analysis
Proposition 1. TIP is invariant under isomorphism.
@@
Complexity. PH can be efficiently computed for dimensions 0 and 1, with a worst-case time complexity of $O(m\alpha(m))$, where $m$ represents the number of sorted edges in a graph. Here, $\alpha(\cdot)$ represents the inverse Ackermann function, which is extremely slow-growing and can essentially be considered as a constant for practical purposes. Therefore, the primary factor that affects the calculation of PH is the complexity of sorting all the edges, which is $O(m\log m)$. Our resampling and persistence encoding mechanism ensures that the coarsened graphs are sparse rather than dense, making our approach both efficient and scalable. We provide running time comparisons in Appendix E.2, which indicates that the inclusion of TIP does not impose a significant computational burden.
@@ 5.1 Experimental Setup
To investigate the effectiveness of PH in GP, we integrate TIP with DiffPool, MinCutPool, and DMoNPool, which are the pioneering approaches that have inspired many other pooling methods. Additionally, as most pooling methods rely on GNNs as their backbone, we compare the widely used GNN models GCN [24], GIN [44], and GraphSage [15]. We also look into another two related and State-of-the-Art GNN models, namely TOGL [18] and GSN [4], which incorporate topological information and graph substructures into GNNs to enhance the expressive power. Several other GP methods, namely Graclus [7] and TopK [11] are also compared. For model selection, we follow the guidelines provided by the original authors or benchmarking papers. Our method acts as an additional plug-in to existing pooling methods (referred to as -TIP) without modifying the remaining model structure and hyperparameters. Appendix B.1 provides detailed configurations of these models.
@@
To evaluate the capabilities of our model across diverse domains, we assess its performance on a variety of graph datasets commonly used in graph classification tasks. We select several benchmarks from TU datasets [30] and OGB datasets [19]. Specifically, we adopt molecular datasets NCI1, NCI109, and OGBG-MOLHIV, bioinformatics datasets ENZYMES, PROTEINS, and DD, as well as social network datasets IMDB-BINARY and IMDB-MULTI. Furthermore, to investigate the topology-preserving ability of our method, we conduct experiments on several highly structured datasets (ring, torus, grid2d) obtained from the PyGSP library. Appendix B.2 provides detailed statistics of the datasets.
@@ 5.2 Preserving Topological Structure
In this experiment, we study Q1 about the ability of PH to preserve topological structure during pooling. Specifically, we assess the topological similarity between the original and coarsened graphs $\mathcal{G}$ and $\mathcal{G^{\prime}}$, by comparing the Wasserstein distance associated with their respective PDs $\mathcal{D}_{1}$ and $\mathcal{D^{\prime}}_{1}$. This evaluation criterion is widely used to compare the topological similarity of graphs [45, 39]. We utilize Forman curvature on each edge of the graph as the filtration, which incorporates edge weights and graph clusters to better capture the topological features of the coarsened graphs [40, 42]. We consider the 1-Wasserstein distance $W\left(\mathcal{D}_{1},\mathcal{D}^{\prime}_{1}\right)=\inf_{\gamma\in\Pi\left(\mathcal{D}_{1},\mathcal{D}^{\prime}_{1}\right)}\mathbb{E}_{(x,y)\sim\gamma}[\|x-y\|]$ as the evaluation metric, where $\Pi(\cdot)$ is the set of joint distributions $\gamma(x,y)$ whose marginals are $\mathcal{D}_{1}$ and $\mathcal{D}^{\prime}_{1}$, respectively. Note that we are not learning a new filtration but keep a fixed one. Rather, we use learnable filtrations in training to enhance flexibility, and solely optimize $L_{topo}$ as the main objective.
Table 1: Results to show the topology-preserving ability. Wasserstein distance (lower is better) is used to assess the topological similarity. A bold value indicates the overall winner.
@@
Table 2: Test accuracy (higher is better) of graph classification on benchmark datasets. A bold value indicates the overall winner. Gray background indicates that TIP outperforms the base GP.
@@ 5.3 Graph Classification
In this experiment, we examine the impact of PH on GP in downstream tasks to answer Q2. We have observed in the former experiment that PH can preserve essential topological information during pooling. However, two additional concerns arise: (1) Does TIP continue to generate invariant sub-topology in the downstream task? (2) If so, does this sub-topology contribute to the performance of the downstream task? To address these concerns, we evaluate TIP using various graph classification benchmarks, where the accuracy achieved on these benchmarks serves as a measure of a method’s ability to selectively preserve crucial information based on the task at hand.
We begin by visualizing the coarsened graphs in this task, where edges are cut-off by a small value. From Fig. 4, we can clearly observe that our method manage to preserve the essential sub-topology similar to the original graphs, while dense pooling methods cannot preserve any topology. As discussed in [29], dense pooling methods achieve comparable performance when the assignment matrix $\mathbf{S}$ is replaced by a random matrix. Here our visualization reveals that regardless of the value of $\mathbf{S}$, the coarsened graph always approaches to a fully connected one. Sparse pooling methods, on the other hand, manage to preserve some local structures through clustering or dropping, but the essential global topological structures are destroyed.
Table 2 presents the average and standard deviation of the graph classification accuracy on benchmark datasets, where the results of GP and several baseline GNNs are provided. Experimental results demonstrate that TIP can consistently enhance the performance of the three dense pooling methods. While the original dense pooling methods sometimes underperform compared to the baselines, they are able to surpass them after integrating TIP.
@@
Moreover, an intriguing observation can be found on ENZYMES dataset, where TOGL surpasses the baseline GNNs. TOGL in practice, incorporates PH into GNNs (GraphSage in our implementation), so this results underscores the significance of incorporating topological information for improved performance on ENZYMES. Further, our method demonstrates more significant improvements by augmenting the three dense pooling methods on the ENZYMES dataset. One possible explanation for the observed phenomenon is that the coarsened graphs generated by our methods bear a striking resemblance to numerous frequent subgraphs present in this dataset [10]. Such substructures may serve as indicators of unique characteristics within the graph, rendering them valuable for subsequent tasks. However, it is also worth noting that TOGL only exhibits marginal improvements or even underperforms on the other datasets. This suggests that simply integrating PH features into GNN layers does not fully exploit topological information. Conversely, injecting global topological invariance into pooling layers in our method yields superior performance.
To demonstrate the effectiveness of preserving the invariant sub-topology, we compared DiffPool-TIP with its variant counterpart, DiffPool-TIP-NL (no topological loss), by replacing $\mathcal{L}_{topo}$ with the original $\mathcal{L}_{r}$ in DiffPool (see Table 3 in Appendix A). The training objective curve and the Wasserstein distance curve are presented in Figure 5, both based on the ENZYMES dataset and a fixed filtration (the same as in Section 5.2). From the figures, it is evident that the objective value decreases as the coarsened graphs become more similar in topology to the original graphs when using DiffPool-TIP. However, when training without $\mathcal{L}_{topo}$, the performance is inferior. Additionally, even when the objective value converges, DiffPool-TIP-NL still exhibits changing topology, whereas DiffPool-TIP maintains a stable topology, possibly benefiting from the stability of PH [39]. This also suggests that multiple suboptimal topologies may contribute equally to the objective. Our topology invariant pooling strategy consistently selects topologies similar to the original graph, which leads to better performance. Additional visualization results and analysis about the coarsened graphs obtained by DiffPool-TIP-NL can be found in Appendix E.3.
Besides, we design an additional set of experiments in Appendix E.4, where the topological structure of the graph is highly task-relevant. Ablation study about the contributions of different modules are shown in Appendix E.5. Finally, to empirically demonstrate the expressive power of our proposed method, we provide an experiment on distinguishing non-isomorphic graphs in Appendix E.6.
Figure 5: The training curves of DiffPool-TIP and DiffPool-TIP-NL on ENZYMES dataset. We show the average values and min-max range of objective and Wasserstein distance for multiple runs.
@@ B.1 Implementation detail
In the graph classification task, all datasets are splitted into train (80%), validation (10%), and test (10%) data. Following the evaluation protocol in [47, 29], we train all models using the Adam optimizer [23] and implement a learning rate decay mechanism, reducing the learning rate from $10^{-3}$ to $10^{-5}$ with a decay ratio of 0.5 and a patience of 10 epochs. Additionally, we use early stopping based on the validation accuracy with patience of 50 epochs. We report statistics of the performance metrics over 20 runs with different seeds.
@@
All the methods are implemented using PyTorch and PyG [35, 9]. The compared methods are implemented following the implementations provided in the PyG library  [^1] https://pytorch-geometric.readthedocs.io/en/latest/modules/nn.html. In the case of DiffPool, it uses a 3-layer GraphSage in each pooling layer, while MinCutPool and DMoNPool use a 1-layer GCN before pooling and a 1-layer GNN [31] in each pooling layer. Note that in DiffPool, the GNNs in Eqs. 2 and 4 are different, while in MinCut and DNoNPool they are the same one, as what their do in the original papers. TopK and Graclus are based on a 1-layer GNN [31]. TOGL is implemented using a 3-layer GraphSage as it has demonstrated superior performance on graph classification tasks (see Table 2). For the baseline GNN models (GCN, GIN, and GraphSage), we use 3 layers with mean/max pooling. In our model, TIP is incorporated as a plugin to existing pooling methods, without modifying the remaining model structure and hyperparameters. We replace the reconstruction loss $\mathcal{L}_{r}$ with $\mathcal{L}_{topo}$ while keeping the cluster loss $\mathcal{L}_{c}$ unchanged. In the case of MinCutPool and DMoNPool, our resampling strategy is added after their original normalization of the coarsened graphs. In preserving topological structure experiments, we initialize node features as the concatenation of the first ten eigenvectors of graph Laplacian matrices. Moreover, we follow the settings in previous works [18, 17] to extend $\mathcal{D}_{1}$ as follows: (1) each cycle is paired with the edge that created it; (2) edges $e$ that do not create a cycle (still in this circle) are assigned a ‘dummy’ tuple value, such as $(f(e),f(e))$; (3) all other edges will be paired with the maximum value of the filtration $f_{max}$. In practice we set $f_{max}$ plus a constant as infinity of destruction time. Therefore, $\mathcal{D}_{1}$ consists of as many tuples as the number of edges $m$.
@@ Appendix C Theoretical Expressivity of TIP
Proof.
Assume that $\mathcal{G}$ and $\mathcal{G}^{\prime}$ have $n$ and $n^{\prime}$ nodes, and the label sequences of them diverge at some iteration $h$, which means there exists at least one label whose count is unique. Let nodes $u$ and $u^{\prime}$ be the nodes with unique count in $\mathcal{G}$ and $\mathcal{G}^{\prime}$, respectively. Denote $La^{(h)}:=\{l_{1},l_{2},...\}$ as an enumeration of the finitely many hashed labels at iteration $h$. We can build a filtration function $f$ by assigning a vertex $v$ with label $l_{i}$ to its index, i.e. $f(v):=i$ except that $f(u)=n+n^{\prime}+1$ and $f(u^{\prime})=n+n^{\prime}+2$. The filtration of edge $(u,w)$ is defined as $f(v,w):=max\{f(v),f(w)\}$, and for isolated nodes $v$, the filtration of self-loop edges is $f(v,v)=f(v)$. Therefore, node with unique label count and its connected edges always correspond to the largest filtration value. Note that the 1-dimensional PD has been extended to have the same cardinality as the number of edges. If node $u$ or $u^{\prime}$ forms a circle, the creation of this circle is related to the edge with the largest filtration; if node $u$ or $u^{\prime}$ does not form a circle, the corresponding edges lie on the diagonal of $\mathcal{D}_{1}$ with unique coordinates; otherwise node $u$ constitute a circle while $u^{\prime}$ does not, then the corresponding edges lie in different parts in $\mathcal{D}_{1}$ and $\mathcal{D}^{\prime}_{1}$. Hence, $\mathcal{D}_{1}\neq\mathcal{D}^{\prime}_{1}$.
@@
To demonstrate that TIP is more expressive than other dense pooling methods, we provide examples of graph pairs that cannot be distinguished by 1-WL but can be by TIP. We present an example of such non-isomorphic graphs in Fig. 6, where in the second graph the edge connecting two triangles does not form a circle. This edge corresponds to zero persistence and is eliminated in TIP. Consequently, the two originally non-isomorphic graphs can be easily distinguished. Provided that the three sufficient conditions proposed in [3] are satisfied, the pooling layers retain the same level of expressive power as GNN. In TIP, the reduction of node features remains unaltered, thereby fulfilling the three conditions. Additionally, TIP is capable of distinguishing certain non-isomorphic graphs, indicating its superior expressive power compared to conventional dense pooling methods such as DiffPool, MinCutPool, and DMoNPool.
Proposition 1. TIP is invariant under isomorphism.
@@
Lemma 1 (Theorem 2 in [37]). Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be two isomorphic graphs. For any equivariant filtration $f$, the corresponding persistence diagrams are equal.
@@
Proof.
Let $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ be two isomorphic graphs with $n$ nodes. Their node features after applying several layers of GNNs are denoted as $\mathbf{X}_{1}=\{\mathbf{x}_{i}\}_{i=1}^{n}$ and $\mathbf{X}_{2}=\{\mathbf{y}_{i}\}_{i=1}^{n}$, respectively. Let the assignment matrix of $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$ obtained in Eq. 2 be $\mathbf{S}_{1}=\left\{\left\{s_{ij}\right\}_{i=1}^{n}\right\}_{j=1}^{n^{\prime}}$ and $\mathbf{S}_{2}=\left\{\left\{q_{ij}\right\}_{i=1}^{n}\right\}_{j=1}^{n^{\prime}}$ with ${\textstyle\sum_{j=1}^{n^{\prime}}}s_{ij}=1$ and ${\textstyle\sum_{j=1}^{n^{\prime}}}q_{ij}=1$. Let the coarsened graphs after pooling as $\mathcal{G}_{1P}$ and $\mathcal{G}_{2P}$ and with $n^{\prime}$ nodes, and their corresponding node features as $\mathbf{X}_{P1}=\{\mathbf{x}_{Pj}\}_{j=1}^{n^{\prime}}$ and $\mathbf{X}_{P2}=\{\mathbf{y}_{Pj}\}_{j=1}^{n^{\prime}}$, where $x_{Pj}=\textstyle\sum_{i=1}^{n}\mathbf{x}_{i}\cdot\mathbf{s}_{ij}$ and $y_{Pj}=\textstyle\sum_{i=1}^{n}\mathbf{y}_{i}\cdot\mathbf{q}_{ij}$ according to Eq. 4.
@@
which indicates that the output of TIP is invariant given isomorphic graphs as input. We have already proved that isomorphic graph pairs after TIP are still isomorphic, so Eq. 9 still holds for multiple layers of pooling. ∎